& A hot-air balloon must support its own weight plus a person for a total weight of 1200 N Ambient air is at 20C and 1 atm . The hot air inside the balloon is at 65C and 1 atm. What diameter of spherical balloon just will support the total weight ? neglect the size of the hot-air inlet vent .
1
Expert's answer
2011-06-20T13:30:37-0400
Archimedes' theorem: the ball will support the total weight in the air, if the density of the ball with air + cargo equal to the density of air at 20 degrees.

1200 N = 120 kg
ρsys = (120 +V* ρ 65)/V = ρ20
ρ65 = 1.045 kg/m3 - density at 65 degrees
ρ20 = 1.204 kg/m3 - density at 20 degrees

thus
V = 755 m3 => r = (3V/ 4 π)1/3 = (180) 1/3 = 5.65 m
Hence diameter D = 2r = 11.3m